# LLM DNA: Tracing Model Evolution via Functional Representations

### v1

@@ Abstract
The explosive growth of large language models (LLMs) has created a vast but opaque landscape: millions of models exist, yet their evolutionary relationships through fine-tuning, distillation, or adaptation are often undocumented or unclear, complicating LLM management. Existing methods are limited by task specificity, fixed model sets, or strict assumptions about tokenizers or architectures. Inspired by biological DNA, we address these limitations by mathematically defining LLM DNA as a low-dimensional, bi-Lipschitz representation of functional behavior. We prove that LLM DNA satisfies inheritance and genetic determinism properties and establish the existence of DNA. Building on this theory, we derive a general, scalable, training-free pipeline for DNA extraction. In experiments across 305 LLMs, DNA aligns with prior studies on limited subsets and achieves superior or competitive performance on specific tasks. Beyond these tasks, DNA comparisons uncover previously undocumented relationships among LLMs. We further construct the evolutionary tree of LLMs using phylogenetic algorithms, which align with shifts from encoder-decoder to decoder-only architectures, reflect temporal progression, and reveal distinct evolutionary speeds across LLM families.
@@ 1 Introduction
Existing approaches to defining and extracting LLM “DNA” remain premature. Most prior work compresses LLMs into task-specific representations for model routing (Ong et al., 2025) or ensemble learning (Huang et al., 2024; Fang et al., 2024). These representations are trained for specific objectives and neither generalize across tasks nor capture overall model characteristics. Zhuang et al. (2024) propose EmbedLLM, the first task-agnostic approach, which learns a compact representation for multiple downstream tasks. However, this representation is defined relative to a fixed set of LLMs—it changes when new models are added—and thus is not an intrinsic DNA of each model. Nikolic et al. (2025) and Zhu et al. (2025) further explore how to measure provenance or independence between two LLMs by calculating token similarity or parameter correlation, but these methods do not generalize to diverse LLMs with varying architectures and tokenizers. Collectively, these approaches do not furnish a formal and general definition of LLM DNA or a practical method for its extraction.
@@
In this paper, we formally define LLM DNA as a low-dimensional vector that obtained via a bi-Lipschitz map from the LLM functional space. We prove that this definition naturally satisfies inheritance and genetic determination and, using the Johnson–Lindenstrauss (JL) lemma (Johnson et al., 1984), formally establish the existence of such DNA. To practically extract LLM DNA, we derive random linear projection as an effective extraction method from these theoretical foundations and design a general, scalable, training-free pipeline for DNA extraction. Based on DNAs, we build the evolutionary tree for 305 LLMs using established algorithms and report several new findings.
Contributions. (1) We formally define LLM DNA and prove that it satisfies inheritance and genetic determination properties, proving the existence of LLM DNA. (2) Building on these results, we propose a general, scalable, training-free pipeline for extracting LLM DNA. (3) Experiments on 305 models show that LLM DNA outperforms existing baselines on specific tasks and aligns with prior findings. (4) Using LLM DNAs, we construct a phylogenetic tree that reveals undocumented relationships among LLMs, aligns with known architectural and temporal shifts, and highlights differing evolutionary rates across LLM families.
@@ 2 Related Work
LLM DNA relates to prior work using terms such as representation and fingerprint. We group this literature into two views: (i) a macroscopic view, which manages large groups of models by mapping each model to a representation for model routing and ensemble learning.; and (ii) a microscopic view, which uses fingerprints to assess whether two LLMs are related for provenance or identity.
Representation of LLMs. This research direction manages large LLM collections by encoding each model into a representation. Most methods are task-specific: Ong et al. (2025); Wang et al. (2024) learn representations for model routing, while Huang et al. (2024); Fang et al. (2024) encode models for ensemble learning. Such representations are hard to reuse across tasks, and their generalizability is unclear. Zhuang et al. (2024) propose EmbedLLM, a compact embedding supporting multiple tasks (e.g., model routing and benchmark prediction), but it is defined over a fixed model set; adding models requires retraining and can shift the learned embeddings. In contrast, our proposed LLM DNA is a stable intrinsic property extractable independently for each model.
@@
Fingerprint of LLMs. This line of work evaluates dependence or similarity between pairs of LLMs by mapping each model to a representation, or fingerprint. Unlike watermarks, which actively insert a “fingerprint” during training (Zhao et al., 2024; Liu et al., 2024; Tang et al., 2025), LLM fingerprinting does not modify training; it analyzes identifiable properties of the model. For example, several studies test whether a remote LLM API serves the claimed model using strategic queries (Pasquini et al., 2025) or specialized prompts (Tsai et al., 2025) followed by a response classification; Fu et al. (2025) further train a classifier to capture unique stylistic features. Beyond text responses, Nikolic et al. (2025) and Zhu et al. (2025) use parameters and token embeddings to improve provenance prediction. However, unlike work on LLM representations, these pairwise comparisons are highly structure-dependent and remain limited for large-scale management of heterogeneous LLMs. We propose LLM DNA to bridge this gap by introducing a general representation that captures structural relationships among heterogeneous LLMs.
3 Definition of LLM DNA
This section first motivates our study by outlining how LLM research parallels biological research, followed by a formal definition of LLM DNA in Section 3.1. Furthermore, Section 3.2 elaborates that this definition satisfies the desired properties of biological DNA.
Biological Analogy and Motivation. The 1953 discovery of DNA, followed by sequencing and editing technologies, fundamentally transformed biological research. DNA has two key properties: (i) genetic determinism—the sequence specifies an organism’s characteristics; and (ii) inheritance—transmission across generations. These properties allow researchers to infer species’ relations by comparing DNA and to introduce traits by appending designed sequences. This motivates asking whether LLMs possess a DNA analogue with comparable properties; if so, how it can be extracted and used. Formalizing LLM DNA would enable methods inspired by molecular biology to advance LLM research.
@@ 3.1 Definition of LLM DNA
We formally define an LLM as a function from a discrete sequence of input texts to a vector of continuous logits.
@@
This definition applies to general LLMs with diverse architectures and adopts an end-to-end perspective. For theoretical analysis, we assume the input length is bounded by a very large constant $m$; however, our practical stochastic-approximation pipeline in Section 5 does not depend on any particular $m$. Under this bound, the input space becomes finite, which directly yields Lemma 3.1, a key step in establishing the existence of LLM DNA.
{restatable}
[LLM Functional Space is a Hilbert Space]lemmalemishilbert Let $\mathcal{F}$ be the space of LLM functions mapping from the finite input space $\mathcal{S}_{m}=\{x_{1},\dots,x_{N}\}$ to the vector space of outputs $\mathcal{O}:=\mathbb{R}^{N}$. Let $\mu:\mathcal{S}_{m}\to[0,1]$ be a probability distribution on the input space, assigning a probability $\mu(x_{i})$ to each sequence such that $\sum_{i=1}^{N}\mu(x_{i})=1$. The pair $(\mathcal{F},d_{H})$ forms a Hilbert space, where the distance metric $d_{H}$ is defined as: $d_{H}(f_{1},f_{2})=\sqrt{\sum_{i=1}^{N}\mu(x_{i})\|f_{1}(x_{i})-f_{2}(x_{i})\|_{2}^{2}}$
The proof of Lemma 3.1 appears in Appendix A.1. With the LLM formally defined, we further define the DNA of an LLM, a low-dimensional representation that preserves functional distance.
@@
Definition 3.2 guarantees that functionally similar LLMs have proximate DNAs, and vice versa. It is crucial to note that this bi-Lipschitz condition is imposed on the mapping from the function space $\mathcal{F}$ to the DNA space $\mathcal{D}$. This property governs the geometry between the two spaces, ensuring that distances between different functions ($f_{1}$ and $f_{2}$) are faithfully preserved. Notably, the existence of a bi-Lipschitz DNA representation neither relies on nor implies that any individual LLM $f:\mathcal{S}_{m}\to\mathcal{O}$ must itself be a bi-Lipschitz function.
3.2 Properties of LLM DNA
@@
[Inheritance]theoremthminheritance Let an LLM $f_{2}\in\mathcal{F}$ be derived from an LLM $f_{1}\in\mathcal{F}$, with corresponding DNAs $\tau_{f_{2}}$ and $\tau_{f_{1}}$. For any desired DNA proximity $\epsilon_{\tau}>0$, there exists a Hilbert distance threshold $\delta_{H}>0$ such that if $d_{H}(f_{1},f_{2})<\delta_{H}$, then $d_{\tau}(\tau_{f_{1}},\tau_{f_{2}})<\epsilon_{\tau}$.
@@
[Genetic Determinism]theoremthmdeterminism Let $f_{1},f_{2}\in\mathcal{F}$ be two LLMs with corresponding DNAs $\tau_{f_{1}},\tau_{f_{2}}\in\mathcal{D}$. For any desired functional proximity $\epsilon_{H}>0$, there exists a DNA distance threshold $\delta_{\tau}>0$ such that if $d_{\tau}(\tau_{f_{1}},\tau_{f_{2}})<\delta_{\tau}$, then $d_{H}(f_{1},f_{2})<\epsilon_{H}$.
The proofs of Theorem 3.2 and Theorem 3.2 appear in Appendix A.2 and Appendix A.3, respectively. Both properties are direct consequences of the lower and upper bounds implied by the bi-Lipschitz constraint. This connection suggests that LLM DNA can be used to study LLM evolutions with methodologies analogous to those employed in biology.
4 Existence of LLM DNA
This section provides a formal proof for the existence of the LLM DNA. The core idea is that an LLM’s function can be represented as a vector in a high-dimensional Hilbert space (Lemma 3.1). This allows us to apply the Johnson-Lindenstrauss (JL) Lemma (Johnson et al., 1984), a powerful result in dimensionality reduction, to guarantee the existence of a low-dimensional embedding that preserves the geometry of the LLM functional space $\mathcal{F}$.
{restatable}
[Existence of LLM DNA]theoremthmdnaexistence For any finite set of $K$ LLMs $\mathcal{F}_{K}=\{f_{1},\dots,f_{K}\}\subset\mathcal{F}$, a DNA representation (Definition 3.2) with DNA space $\mathcal{D}=\mathbb{R}^{L}$ exists, satisfying for all $f_{i},f_{j}\in\mathcal{F}_{K}$, $c_{1}\cdot d_{H}(f_{i},f_{j})\leq d_{\tau}(\tau_{f_{i}},\tau_{f_{j}})\leq c_{2}\cdot d_{H}(f_{i},f_{j}),$ where the target dimension $L$ is $L=O\left(\left(\frac{c_{2}+c_{1}}{c_{2}-c_{1}}\right)^{2}\log K\right).$
The proof of Theorem 4 appears in Appendix A.4. The required DNA dimension $L$ trades off with the bi-Lipschitz constants $c_{1},c_{2}$, whose tightness—captured by $\epsilon=(c_{2}-c_{1})/(c_{2}+c_{1})$—governs embedding distortion. High fidelity (small distortion) requires $c_{1}\approx c_{2}$, yielding very small $\epsilon$ and, by $L=O(\epsilon^{-2}\log K)$, a large $L$. Conversely, tolerating more distortion (a larger gap between $c_{1}$ and $c_{2}$) increases $\epsilon$ and permits a more compact, lower-dimensional DNA representation.
Theorem 4 not only guarantees the existence of an LLM DNA but also yields a practical construction. The proof leverages the constructive nature of the JL lemma: the required linear map $E$ need not be specially engineered; a scaled random linear projection satisfies the desired bi-Lipschitz property with arbitrarily high probability. This insight is formalized in the following corollary.
Corollary 4.1 (Construct LLM DNA via Random Projection).
For any finite set of $K$ LLMs $\mathcal{F}_{K}=\{f_{1},\dots,f_{K}\}\subset\mathcal{F}$, the LLM DNA in Theorem 4 can be obtained from a random linear projection $E:\mathcal{F}\to\mathbb{R}^{L}$ with probability at least $1-2/K^{2}$.
Proof.
Since Lemma 3.1 shows that $\mathcal{F}$ is a Hilbert space, the JL lemma applies. Invoking a standard formulation of the lemma (e.g., Theorem 2.1 in Dasgupta & Gupta (2003)) for the set of vector representations of LLMs in $\mathcal{F}_{K}$ yields the desired projection and associated probability bound. ∎
@@ 5 Extract LLM DNA
Extracting the LLM DNA in practice poses two main challenges. The first is to instantiate the abstract output space $\mathcal{O}$ in a way that captures semantic meaning, since direct string-level comparison of LLM outputs is inadequate. The second is computational infeasibility: the theoretical Hilbert distance $d_{H}$ requires summing output differences over the entire, combinatorially large input space $\mathcal{S}_{m}$. We address these challenges by introducing a semantic-aware representation of responses (Section 5.1) and by defining an expectation-based functional distance that admits tractable approximation (Section 5.2). The overall pipeline is further elaborated in Section 5.3. The complete procedure is visualized in Figure 1.
Figure 1: Visualization of LLM DNA extraction workflow.
@@ 5.1 Construct Semantic-Aware Representation
This subsection details how to build a semantic-aware vector representation $\phi(f(x_{i}))$ for each LLM response $x_{i}$ such that $\|\phi(f_{1}(x_{i}))-\phi(f_{2}(x_{i}))\|_{2}^{2}$ effectively approximates $\|f_{1}(x_{j})-f_{2}(x_{j})\|_{2}^{2}$. Purely string-based comparisons ignore semantic information: for example, the words “holiday” and “vacation” are semantically similar yet differ completely as character sequences. To convert textual outputs into vectors that capture semantic meaning, we use a sentence-embedding model to map each response to a fixed-size embedding.
This design offers three advantages: (1) it is LLM-agnostic across text-generation models, avoiding dependencies on internal architectures or tokenizers required by many embedding-based or model-based extraction methods; (2) it supports evaluation of closed-source LLMs, since API-only models can still be queried to obtain responses for DNA computation; and (3) it extends seamlessly to new LLMs without altering previously computed DNAs, thereby addressing the limitation noted in EmbedLLM (Zhuang et al., 2024).
@@ 5.2 Approximate LLM Functional Distance
This subsection details our practical approach to measuring the distance between LLMs. The Hilbert distance $d_{H}$ introduced in Lemma 3.1 provides a complete theoretical metric, but its computation requires evaluating functions over the entire, combinatorially large input space $\mathcal{S}_{m}$, which is infeasible. To bridge theory and practice, we introduce a tractable alternative: a Stochastic Functional Distance. This new metric is defined not over the full input space, but as a statistical expectation over a random sample of inputs, reflecting the practical scenario of evaluating models on a finite set of representative prompts.

### v2

@@ Abstract
The explosive growth of large language models (LLMs) has created a vast but opaque landscape: millions of models exist, yet their evolutionary relationships through fine-tuning, distillation, or adaptation are often undocumented or unclear, complicating LLM management. Existing methods are limited by task specificity, fixed model sets, or strict assumptions about tokenizers or architectures. Inspired by biological DNA, we address these limitations by mathematically defining LLM DNA as a low-dimensional, bi-Lipschitz representation of functional behavior. We prove that LLM DNA satisfies inheritance and genetic determinism properties and establish the existence of DNA. Building on this theory, we derive a general, scalable, training-free pipeline for DNA extraction. In experiments across 305 LLMs, DNA aligns with prior studies on limited subsets and achieves superior or competitive performance on various tasks. Beyond these tasks, DNA comparisons uncover previously undocumented relationships among LLMs. We further construct the evolutionary tree of LLMs using phylogenetic algorithms, which align with shifts from encoder-decoder to decoder-only architectures, reflect temporal progression, and reveal distinct evolutionary speeds across LLM families. dna.xtra.science github.com/Xtra-Computing/LLM-DNA pypi.org/project/llm-dna
@@ 1 Introduction
Existing approaches to defining and extracting LLM “DNA” remain premature. Most prior work compresses LLMs into task-specific representations for model routing (Ong et al., 2025) or ensemble learning (Huang et al., 2024). These representations are trained for specific objectives and neither generalize across tasks nor capture overall model characteristics. Zhuang et al. (2025) propose EmbedLLM, the first task-agnostic approach, which learns a compact representation for multiple downstream tasks. However, this representation is defined relative to a fixed set of LLMs—it changes when new models are added—and thus is not an intrinsic DNA of each model. Nikolic et al. (2025) and Zhu et al. (2025) further explore how to measure provenance or independence between two LLMs by calculating token similarity or parameter correlation, but these methods do not generalize to diverse LLMs with varying architectures and tokenizers. Collectively, these approaches do not furnish a formal and general definition of LLM DNA or a practical method for its extraction.
@@
In this paper, we formally define LLM DNA as a low-dimensional vector that obtained via a bi-Lipschitz map from the LLM functional space. We prove that this definition naturally satisfies inheritance and genetic determination and, using the Johnson–Lindenstrauss (JL) lemma (Johnson et al., 1984), formally establish the existence of such DNA. To practically extract LLM DNA, we derive random linear projection as an effective extraction method from these theoretical foundations and design RepTrace—a general, scalable, training-free pipeline for DNA extraction. Finally, we build the evolutionary tree for 305 LLMs using RepTrace and report various new findings.
Contributions. (1) We formally define LLM DNA and prove that it satisfies inheritance and genetic determination properties, proving the existence of LLM DNA. (2) Building on these results, we propose a general, scalable, training-free pipeline for extracting LLM DNA. (3) Experiments on 305 models show that LLM DNA outperforms existing baselines on specific tasks and aligns with prior findings. (4) Using LLM DNAs, we construct a phylogenetic tree that aligns with known architectural and temporal shifts and reveals undocumented relationships among LLMs.
@@ 2 Related Work
LLM DNA relates to prior work using terms such as representation and fingerprint. We group this literature into two views: (i) a macroscopic view, which manages large groups of models by mapping each model to a representation for model routing and ensemble learning; and (ii) a microscopic view, which uses fingerprints to assess whether two LLMs are related for provenance or identity.
Representation of LLMs. This direction manages large LLM collections by encoding each LLM into a representation. Most methods are task-specific: Ong et al. (2025); Wang et al. (2024) learn representations for model routing, while Huang et al. (2024); Fang et al. (2024) encode models for ensemble learning. For instance, HybridLLM (Ding et al., 2024), trains a binary “easy-hard” classifier to route queries to a small or a larger model. RouteLLM (Ong et al., 2025) learns a supervised router that selects between stronger and weaker models based on query features. MetaLLM (Nguyen et al., 2024) formulates routing as a multi-armed bandit, adaptively trading off expected answer quality against invocation cost via online feedback. LLM-ensemble (Fang et al., 2024) learns the weights for different LLMs to predict a specific attribute value. These task-specific approaches rely on supervised training for particular tasks, making their learned representations not generalizable across tasks. Zhuang et al. (2025) propose EmbedLLM, the first compact embedding supporting multiple tasks (e.g., model routing and benchmark prediction), but it is defined over a fixed model set; adding models requires retraining and can shift the learned embeddings. Yax et al. (2025) propose PhyloLM, a state-of-the-art approach for constructing LLM phylogenies. It defines inter-model distances in a tokenizer-aware manner: for models with the same tokenizer, it uses Nei’s genetic distance over output token distributions; for different tokenizers, it approximates distance using the first four characters of concatenated tokens. As a result, the cross-tokenizer metric is dominated by early-token behavior and may not reflect sentence-level distributions. In contrast, our LLM DNA is a stable and generalizable sentence-level intrinsic property extractable independently for each LLM.
@@
Fingerprint of LLMs. This line of work evaluates dependence or similarity between pairs of LLMs by mapping each model to a representation, or fingerprint. Unlike watermarks, which actively insert a fingerprint during training (Zhao et al., 2024; Liu et al., 2024; Tang et al., 2025), fingerprinting is post hoc and analyzes identifiable model properties without modifying training. Prior studies verify whether a remote LLM API serves the claimed model using strategic queries (Pasquini et al., 2025) or specialized prompts (Tsai et al., 2025) with response classification; Fu et al. (2025) further trains classifiers to capture stylistic signatures. Others go beyond text, leveraging parameters or token embeddings for provenance prediction (Nikolic et al., 2025; Zhu et al., 2025). However, these pairwise methods are often structure-dependent and hard to scale to heterogeneous LLMs. LLM DNA, however, is a general representation capturing structural relationships across diverse models.
3 Definition and Existence of LLM DNA
This section provides a formal definition of LLM DNA in Section 3.1. Furthermore, Section 3.2 proves the existence of DNA under this definition. Additionally, Appendix A.1 elaborates that this definition satisfies two properties analogous to those of biological DNA.
@@ 3.1 Definition of LLM DNA
We formally define an LLM as a function from a discrete sequence of input texts to a vector of continuous logits (Definition 3.1). This definition applies to general LLMs with diverse architectures and adopts an end-to-end perspective. For the theoretical analysis, we assume the input length is finite, i.e., bounded by a constant $m$. Notably, “finiteness” does not imply smallness; the sets involved may still be arbitrarily large. Meanwhile, our practical stochastic-approximation pipeline in Section 4 does not depend on any particular value of $m$. Under this bound, the input space becomes finite, which directly yields Lemma A.4, a key step in establishing the existence of LLM DNA.
@@
Definition 3.2 guarantees that functionally similar LLMs have proximate DNAs, and vice versa. These lower and upper bounds in the bi-Lipschitz condition naturally guarantee two key properties of LLM DNA: (1) Inheritance—a lineage operation on an LLM, such as fine-tuning, produces descendants with similar DNAs; and (2) Genetic determinism—LLMs with similar DNAs exhibit similar functional behaviors. Both properties are formally stated and proved in Appendix A.1.
3.2 Existence of LLM DNA
This section provides a formal proof for the existence of the LLM DNA. The core idea is that an LLM’s function can be represented as a vector in a high-dimensional Hilbert space (Lemma A.4). This allows us to apply the Johnson-Lindenstrauss (JL) Lemma (Johnson et al., 1984), a powerful result in dimensionality reduction, to guarantee the existence of a low-dimensional embedding that preserves the geometry of the LLM functional space $\mathcal{F}$.
@@ 3.2 Existence of LLM DNA
[Existence of LLM DNA]theoremthmdnaexistence For any finite set of $K$ LLMs $\mathcal{F}_{K}=\{f_{1},\dots,f_{K}\}\subset\mathcal{F}$, a DNA representation (Definition 3.2) with DNA space $\mathcal{D}=\mathbb{R}^{L}$ exists, satisfying for all $f_{i},f_{j}\in\mathcal{F}_{K}$, $c_{1}\cdot d_{H}(f_{i},f_{j})\leq d_{\tau}(\tau_{f_{i}},\tau_{f_{j}})\leq c_{2}\cdot d_{H}(f_{i},f_{j}),$ where the target dimension $L$ is $L=O\left(\left[(c_{2}+c_{1})/(c_{2}-c_{1})\right]^{2}\log K\right).$
The proof of Theorem 3.2 appears in Appendix A.3. The required DNA dimension $L$ trades off with the bi-Lipschitz constants $c_{1},c_{2}$, whose tightness—captured by $\epsilon=(c_{2}-c_{1})/(c_{2}+c_{1})$—governs embedding distortion. High fidelity (small distortion) requires $c_{1}\approx c_{2}$, yielding very small $\epsilon$ and, by $L=O(\epsilon^{-2}\log K)$, a large $L$. Conversely, tolerating more distortion (a larger gap between $c_{1}$ and $c_{2}$) increases $\epsilon$ and permits a more compact, lower-dimensional DNA representation.
Theorem 3.2 not only guarantees the existence of an LLM DNA but also yields a practical construction. The proof leverages the constructive nature of the JL lemma: the required linear map $E$ need not be specially engineered; a scaled random linear projection satisfies the desired bi-Lipschitz property with arbitrarily high probability. This insight is formalized in the following corollary.
@@
[Construct LLM DNA via Random Projection]corollarycordnaconstruct For any finite set of $K$ LLMs $\mathcal{F}_{K}=\{f_{1},\dots,f_{K}\}\subset\mathcal{F}$, the LLM DNA in Theorem 3.2 can be obtained from a random linear projection $E:\mathcal{F}\to\mathbb{R}^{L}$ with probability at least $1-2/K^{2}$.
@@ 4 Extract LLM DNA
Extracting LLM DNA in practice entails two challenges: (i) instantiating the output space $\mathcal{O}$ to reflect semantic content, since surface-form string comparisons are brittle; and (ii) evaluating functional distance over the full input space $\mathcal{S}_{m}$, which is computationally prohibitive. We propose an extraction pipeline, RepTrace, which addresses these via a semantic response representation (Section 4.1) and an expectation-based functional distance that admits efficient approximation (Section 4.2). The full RepTrace pipeline is provided in Section 4.3 and illustrated in Figure 1.
Figure 1: Visualization of RepTrace: LLM DNA extraction workflow
@@ 4.1 Construct Semantic-Aware Representation
This subsection details how to build a semantic-aware vector representation $\phi(f(x_{i}))$ for each LLM response $x_{i}$ such that $\|\phi(f_{1}(x_{i}))-\phi(f_{2}(x_{i}))\|_{2}$ effectively approximates $\|f_{1}(x_{j})-f_{2}(x_{j})\|_{2}$. Purely string-based comparisons ignore semantic information: for example, the words “holiday” and “vacation” are semantically similar yet differ completely as character sequences. To convert textual outputs into vectors that capture semantic meaning, we use a sentence-embedding model to map each response to a fixed-size embedding.
This design has three benefits: (1) it is LLM-agnostic for text-generation models, requiring no access to internal architectures or tokenizers as in many embedding/model-based methods (Yax et al., 2025); (2) it applies to closed-source LLMs because API-only models can still be queried for DNA computation; and (3) it accommodates new LLMs without recomputing existing DNAs, overcoming the limitation of EmbedLLM (Zhuang et al., 2025).
@@ 4.2 Approximate LLM Functional Distance
This subsection details our practical approach to measuring the distance between LLMs. Though direct computing functional distance is theoretically sound, its computation requires evaluating functions over the entire, combinatorially large input space $\mathcal{S}_{m}$, which is infeasible. To bridge theory and practice, we introduce a tractable alternative: a Stochastic Functional Distance. This new metric is defined not over the full input space, but as a statistical expectation over a random sample of inputs, reflecting the practical scenario of evaluating models on a finite set of representative prompts.